# Diffusion2: Dynamic 3D Content Generation via Score Composition of Orthogonal Diffusion Models

### v3

@@ Abstract
Recent advancements in 3D generation are predominantly propelled by improvements in 3D-aware image diffusion models which are pretrained on Internet-scale image data and fine-tuned on massive 3D data, offering the capability of producing highly consistent multi-view images. However, due to the scarcity of synchronized multi-view video data, it is impractical to adapt this paradigm to 4D generation directly. Despite that, the available video and 3D data are adequate for training video and multi-view diffusion models separately that can provide satisfactory dynamic and geometric priors respectively. To take advantage of both, this paper present Diffusion2, a novel framework for dynamic 3D content creation that reconciles the knowledge about geometric consistency and temporal smoothness from these models to directly sample dense multi-view multi-frame images which can be employed to optimize continuous 4D representation. Specifically, we design a simple yet effective denoising strategy via score composition of pretrained video and multi-view diffusion models based on the probability structure of the target image array. Owing to the high parallelism of the proposed image generation process and the efficiency of the modern 4D reconstruction pipeline, our framework can generate 4D content within few minutes. Additionally, our method circumvents the reliance on 4D data, thereby having the potential to benefit from the scaling of the foundation video and multi-view diffusion models. Extensive experiments demonstrate the efficacy of our proposed framework and its ability to flexibly handle various types of prompts.
@@ 1 Introduction
Figure 1: Diffusion2 is designed to generate dynamic content by creating a dense multi-frame multi-view image matrix in a highly parallel denoising diffusion process with the combination of the foundation video diffusion model and multi-view diffusion model. The generated image can be used to construct a full 4D representation by feeding them into an off-the-shelf 4D reconstruction pipeline.
@@
Spurred by the advances from generative image models ho2020denoising ; song2020denoising ; song2020score ; karras2022elucidating ; zhang2023adding , automatic 3D content creation poole2022dreamfusion ; wang2024prolificdreamer ; tang2023dreamgaussian ; hong2023lrm has witnessed remarkable progress in terms of efficiency, fidelity, diversity, and controllability. Coupled with the breakthroughs in 4D representation yang2023deformable ; wu20234d ; duan20244d , these advances further foster substantial development in dynamic 3D content generation singer2023text ; bahmani20234d ; jiang2023consistent4d ; zhao2023animate124 ; ren2023dreamgaussian4d ; gao2024gaussianflow , which holds significant value across a wide range of applications in animation, film, game, and MetaVerse.
@@
Recently, 3D content generation has achieved considerable breakthroughs in efficiency. Some works liu2023zero ; liu2023syncdreamer ; shi2023mvdream ; wang2023imagedream ; tang2024mvdiffusion++ inject stereo knowledge into the image generation model, enabling these 3D-aware image generators to produce consistent multi-view images, thereby effectively stabilizing and accelerating the optimization. Other efforts hong2023lrm ; chen2024v3d ; voleti2024sv3d ; zuo2024videomv ; tang2024lgm attempt to directly generate 3D representations, such as triplane chen2022tensorf or Gaussian splatting kerbl3Dgaussians . However, the efficiency improvement from these works is largely data-driven wu2023omniobject3d ; reizenstein2021common ; yu2023mvimgnet ; deitke2023objaverse . Consequently, it is infeasible to adapt them to 4D generation due to the scarcity of synchronized multi-view video data. Therefore, most existing 4D generation works jiang2023consistent4d ; yin20234dgen ; ren2023dreamgaussian4d still adopt the score distillation sampling (SDS) approach and suffer from slow or unstable optimization.
@@
However, despite the paucity of 4D data, there are vast available monocular video data and static multi-view image data. Existing works have demonstrated that it is feasible to train diffusion-based generative models learning the distribution of these two classes of data independently  voleti2024sv3d ; liu2023syncdreamer ; tang2024mvdiffusion++ ; blattmann2023align ; blattmann2023stable . Considering that video diffusion model stores the prior of motion and temporal smoothness, and multi-view diffusion model has sound knowledge of geometrical consistency, combining the two generative models to generate 4D assets becomes a highly promising and appealing approach.
To leverage both of them, inspired by the previous work bar2023multidiffusion ; lee2023syncdiffusion , we propose a novel 4D generation framework that reconciles the video and multi-view diffusion priors to directly sample multi-frame multi-view image arrays imitating the photographing process of 4D object. To demonstrate how we realize this combination, we assume that such an image matrix $\mathcal{I}$ has a nice structure: the elements $\mathcal{I}_{i,j}$ and $\mathcal{I}_{k,l}$ not in the same row and column ($i\neq k,j\neq l$) are conditionally independent of each other given $\mathcal{I}_{i,l}$ or $\mathcal{I}_{k,j}$. Based on this property, we design a simple yet effective denoising sampling strategy in which the estimated score is just the convex combination of the scores predicted by two foundation diffusion models. Our formulation is easy to adapt to various prompts including reference image, single-view video, and static 3D content. Unlike the existing SDS-based counterparts, Diffusion2 is highly parallel. Combined with the efficient 4D reconstruction pipeline, our method can generate high-fidelity and diverse 4D assets within just a few minutes. Besides, our approach can also potentially benefit from the further advancement of foundation diffusion models videoworldsimulators2024 .
Our contributions are threefold: (i) We present a novel 4D content generation approach that achieves zero-shot generation of realistic multi-view multi-frame image arrays, which can be integrated with an off-the-shelf modern 4D reconstruction pipeline to efficiently create 4D content. (ii) We identify the conditional independence within elements constituting the image arrays. Based on it, we devise a simple yet effective joint denoising strategy that combines video diffusion model and multi-view diffusion model to directly sample multi-view multi-frame images from their natural joint distribution. (iii) Systematic experiments demonstrate that our proposed method can achieve satisfactory results under different types of prompts, including single image, single-view video, and static 3D content.
@@ 2 Related work
3D generation   3D generation aims at creating static 3D content from different type of conditions like text or reference image. Early efforts employed GAN-based approaches gao2022get3d ; schwarz2020graf . Recently, significant breakthroughs have been achieved with diffusion models ho2020denoising . DreamFusion poole2022dreamfusion introduced score distillation sampling (SDS) to unleash the creativity in diffusion models. A series of subsequent works wang2024prolificdreamer ; shi2023mvdream ; wang2023imagedream ; pan2024enhancing ; tang2023dreamgaussian ; yi2023gaussiandreamer continuously address challenges such as multi-face Janus issues and slow optimization. On the other hand, with the development of large scale 3D datasets yu2023mvimgnet ; deitke2023objaverse , many works try to directly generate 3D contents by using diffusion models. Some studies nichol2022point ; jun2023shap ; hong2023lrm ; tang2024lgm ; wang2024crm have explored the direct generation of 3D representations. Another line of research liu2023zero ; liu2023syncdreamer ; long2023wonder3d ; chen2024v3d ; tang2024mvdiffusion++ focuses on generating multi-view images with sufficient 3D consistency to be used for reconstruction. We also adopt this approach of directly generating consistent images for reconstruction. But unlike the above 3D counterparts, there is no large-scale multi-view synchronized video data. Therefore, we opt to combine geometrical consistency priors and video dynamic priors to generate images.
@@
Video generation   Video generation is an active field that has gained increasing popularity. Recent diffusion-based methods have exhibited unprecedented levels of realism, diversity, and controllability. Video LDM blattmann2023align is the pioneer work to apply the latent diffusion framework rombach2022high to video generation. The subsequent work SVD blattmann2023stable followed its architecture and made effective improvements to the training recipe. W.A.L.T gupta2023photorealistic employed a transformer with window attention tailored for spatiotemporal generative modeling to generate high-resolution videos. VDT lu2023vdt introduced the video diffusion transformer and a spatial-temporal mask mechanism to flexibly capture long-distance spatiotemporal context. The recently introduced SORA videoworldsimulators2024 demonstrated a remarkable capability to generate long videos with intuitively physical fidelity. Models trained on large-scale video data can generate videos with realistic dynamics. Besides, video diffusion models can also provide effective prior for 3D generators chen2024v3d ; han2024vfusion3d ; voleti2024sv3d . Therefore, we build our method on this flourishing domain.
@@
4D generation   Animating category-agnostic objects is challenging and has drawn considerable attention from both academia and industry. Unlike 3D generation, 4D generation requires both consistent geometry and realistic dynamics. Recent works on this domain can be categorized based on the input condition. Some of them create 4D content from text or single image. For instance, MAV3D singer2023text first employs SDS in text-to-4D tasks. 4D-fy bahmani20234d hybridizes different diffusion priors during SDS training. But they suffer from extremely slow generation. DreamGaussian4D ren2023dreamgaussian4d adopts deformable 3D Gaussian yang2023deformable as the underlying 4D representation and exports mesh for texture refinement, significantly improving efficiency. Another line of work predicts dynamic objects from a single-view video with largely dictated motion. Consistent4D jiang2023consistent4d proposes to use SDS approach for geometry consistency and frame interpolation loss for temporal continuity. 4DGen yin20234dgen further grounds the 4D content creation with pseudo labels. Efficient4D pan2024fast mimics a photogrammetry-based neural volumetric video reconstruction pipeline by directly generating multi-view videos for reconstruction. Although we share the same SDS-free design, Efficient4D possesses a strong architectural preference for the fundamental diffusion model and is incapable of synthesizing novel dynamics. Compared to previous works, our framework can efficiently generate diverse 4D contents from different input prompts, avoiding the slow and unstable optimization. In addition, our method has the potential to continuously benefit from the scalability of underlying diffusion models.
@@ 3 Method
As depicted in Figure 2, the proposed 4D generation framework Diffusion2 adopts a two-stage pipeline: dense observation synthesis followed by reconstruction. In Section 3.1 (stage-1), we will elaborate on how to generate dense multi-view multi-frame images for reconstruction through a highly parallelizable denoising process by reconciling the pretrained video diffusion prior and multi-view diffusion prior, and why it is feasible. In Section 3.2 (stage-2), we will briefly introduce how to robustly reconstruct 4D content from the images produced in the first stage.
3.1 Image matrix generation
@@
In this stage, our goal is to generate dense multi-frame multi-view images for reconstruction, which can be denoted as a matrix of images:
@@
Here, $\mathrm{x}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{N}$ is a series of images with $N$ frames or $N$ views, $\nabla_{\mathrm{x}}\log p\left(\mathrm{x};\sigma\right)$ is the score function, which can be estimated as $\nabla_{\mathrm{x}}\log p\left(\mathrm{x};\sigma\right)\approx\left(D_{\theta}(\mathrm{x};\sigma)\right)/\sigma^{2}$ karras2022elucidating ; blattmann2023stable , where $D_{\theta}(\mathrm{x};\sigma)$ is a neural network trained via denoising score matching.
We want to extend the above formulation to the sampling of $\mathcal{I}$. The question is, how do we estimate the score of the joint distribution of $V\times F$ images?
For clarity, denote $\mathcal{I}_{-i,j}\triangleq\{I_{i^{\prime},j}|1\leq i^{\prime}\leq V,i^{\prime}\neq i\}$, $\mathcal{I}_{i,-j}\triangleq\{I_{i,j^{\prime}}|1\leq j^{\prime}\leq F,j^{\prime}\neq j\}$ and $\mathcal{I}_{-i,-j}\triangleq\{I_{i^{\prime},j^{\prime}}|1\leq i^{\prime}\leq V,1\leq j^{\prime}\leq F,i^{\prime}\neq i,j^{\prime}\neq j\}$. We first make an assumption on the structure of $p(\mathcal{I})$.

### v4

# Diffusion2: Dynamic 3D Content Generation via Score Composition of Video and Multi-view Diffusion Models

@@ Abstract
Recent advancements in 3D generation are predominantly propelled by improvements in 3D-aware image diffusion models. These models are pretrained on internet-scale image data and fine-tuned on massive 3D data, offering the capability of producing highly consistent multi-view images. However, due to the scarcity of synchronized multi-view video data, it remains challenging to adapt this paradigm to 4D generation directly. Despite that, the available video and 3D data are adequate for training video and multi-view diffusion models separately that can provide satisfactory dynamic and geometric priors respectively. To take advantage of both, this paper presents Diffusion2, a novel framework for dynamic 3D content creation that reconciles the knowledge about geometric consistency and temporal smoothness from these models to directly sample dense multi-view multi-frame images which can be employed to optimize continuous 4D representation. Specifically, we design a simple yet effective denoising strategy via score composition of pretrained video and multi-view diffusion models based on the probability structure of the target image array. To alleviate the potential conflicts between two heterogeneous scores, we further introduce variance-reducing sampling via interpolated steps, facilitating smooth and stable generation. Owing to the high parallelism of the proposed image generation process and the efficiency of the modern 4D reconstruction pipeline, our framework can generate 4D content within few minutes. Notably, our method circumvents the reliance on expensive and hard-to-scale 4D data, thereby having the potential to benefit from the scaling of the foundation video and multi-view diffusion models. Extensive experiments demonstrate the efficacy of our proposed framework in generating highly seamless and consistent 4D assets under various types of conditions.
@@ 1 Introduction
Spurred by the advances from generative image models (Ho et al., 2020; Song et al., 2021a; b; Karras et al., 2022; Zhang et al., 2023), automatic 3D content creation (Poole et al., 2023; Wang et al., 2023; Tang et al., 2024b; Hong et al., 2024) has witnessed remarkable progress in terms of efficiency, fidelity, diversity, and controllability. Coupled with the breakthroughs in 4D representation (Yang et al., 2024; Wu et al., 2024a; Duan et al., 2024), these advances further foster substantial development in dynamic 3D content generation (Singer et al., 2023; Bahmani et al., 2024; Jiang et al., 2024b; Zhao et al., 2023; Ren et al., 2023; Gao et al., 2024), which holds significant value across a wide range of applications in animation, film, game, and MetaVerse.
@@
Recently, 3D content generation has achieved considerable breakthroughs in efficiency. Some works (Liu et al., 2023; 2024; Shi et al., 2023; Wang & Shi, 2023; Tang et al., 2024c) inject stereo knowledge into the image generation model, enabling these 3D-aware image generators to produce consistent multi-view images, thereby effectively stabilizing and accelerating the optimization. Other efforts (Hong et al., 2024; Chen et al., 2024; Voleti et al., 2024; Zuo et al., 2024; Tang et al., 2024a) attempt to directly generate 3D representations, such as triplane (Chen et al., 2022) or Gaussian splatting (Kerbl et al., 2023). However, the efficiency improvement from these works is largely data-driven (Wu et al., 2023; Reizenstein et al., 2021; Yu et al., 2023; Deitke et al., 2023). Consequently, it is infeasible to adapt them to 4D generation due to the scarcity of synchronized multi-view video data. Therefore, most existing 4D generation works (Jiang et al., 2024b; Yin et al., 2023; Ren et al., 2023) still adopt the score distillation sampling (SDS) approach and suffer from slow or unstable optimization.
@@
However, despite the paucity of 4D data, there are vast available monocular video data and static multi-view image data. Existing works have demonstrated that it is feasible to train diffusion-based generative models learning the distribution of these two classes of data independently  (Voleti et al., 2024; Liu et al., 2024; Tang et al., 2024c; Blattmann et al., 2023b; a). Considering that video diffusion model stores the prior of motion and temporal smoothness, and multi-view diffusion model has sound knowledge of geometrical consistency, combining the two types of generative models to generate 4D assets becomes a highly promising and appealing approach.
Figure 1: Diffusion2 is designed to generate dynamic 3D content by creating a dense multi-frame multi-view image matrix in a highly parallel denoising process with the combination of the foundation video diffusion model and multi-view diffusion model. Please refer to our supplementary demo video for more results.
To leverage both of them, we propose a novel 4D generation framework, Diffusion2, which reconciles the video and multi-view diffusion priors to directly sample multi-frame multi-view image arrays imitating the photographing process of 4D object. Instead of unleashing the pre-trained knowledge stored in model parameters through fine-tuning (Xie et al., 2024; Ren et al., 2024), Diffusion2 achieves this goal in a training-free and architecture-agnostic manner. Our key insight is that the score function for the joint distribution of elements in image array can be approximated by the convex combination of scores for each view and frame due to the conditional independence within the image matrix. Therefore, we can directly sample multi-view multi-frame images within the framework of score-based generative models with off-the-shelf score estimators.
In practice, due to the potential divergence between the learned distributions of two foundational models caused by imbalanced training data, conflicts could emerge during the denoising process. Such conflicts are primarily manifested as two distinct modes in the denoising results, which may compromise the completeness and consistency of the generated images with continuously reducing noise levels. However, we found that this issue can be effectively resolved by leveraging the inherent properties of the diffusion framework, without requiring any additional training. Based on this idea, we further propose the variance-reducing sampling (VRS), which can effectively mitigate such conflicts at the cost of negligible additional time consumption by reconciling the distinct modes at a higher noise level. With this technique, our Diffusion2 can produce smooth multi-view videos with high spatial-temporal consistency. Furthermore, we introduce a dynamic and texture-decoupled reconstruction strategy that transforms the generated discrete image matrix into a continuous 4D representation while minimizing the potential detail loss caused by aggressive variance reduction, especially on images that are far from the reference view. Thanks to the highly parallel denoising in the image matrix synthesis stage and the efficiency of the reconstruction stage, our method can generate high-fidelity and diverse 4D assets within just a few minutes.
Our contributions can be summarized as follows: (i) We develop a novel 4D generation framework that can generate highly consistent multi-view videos in a single pass of reverse diffusion process using only existing multi-view and video diffusion models, without relying on any 4D dataset. The core of this framework is to estimate the score function of the image matrix by a convex combination of the scores predicted by orthogonal diffusion models. We identify the conditional independence within elements constituting the image arrays and establish the theoretical soundness of the proposed approaches based on this property. (ii) We propose the VRS for reconciling heterogeneous scores during denoising, which can effectively alleviate potential conflicts and facilitate the generation of seamless results with high consistency across multiple views and frames. In synergy with it, we also propose a dynamic and texture decoupling reconstruction pipeline to better preserve texture details. (iii) Systematic experiments demonstrate that our proposed method achieves impressive results under different types of prompts. (iv) Notably, our work first proves that it is practical to directly sample highly spatial-temporal consistent multi-view videos of dynamic 3D objects within a single-pass of denoising diffusion procedure, while only using the existing multi-view and video diffusion prior without any additional training on expensive and hard-to-scale 4D dataset.
@@ 2 Related work
3D generation   3D generation aims at creating static 3D content from different type of conditions like text or reference image. Early efforts employed GAN-based approaches (Gao et al., 2022; Schwarz et al., 2020). Recently, significant breakthroughs have been achieved with diffusion models (Ho et al., 2020). DreamFusion (Poole et al., 2023) introduced score distillation sampling (SDS) to unleash the creativity in diffusion models. A series of subsequent works (Wang et al., 2023; Shi et al., 2023; Wang & Shi, 2023; Pan et al., 2024a; Tang et al., 2024b; Yi et al., 2024) continuously address challenges such as multi-face Janus issues and slow optimization. On the other hand, with the development of large scale 3D datasets (Yu et al., 2023; Deitke et al., 2023), many works try to directly generate 3D contents by using diffusion models. Some studies (Nichol et al., 2022; Jun & Nichol, 2023; Hong et al., 2024; Tang et al., 2024a; Wang et al., 2024) have explored the direct generation of 3D representations. Another line of research (Liu et al., 2023; 2024; Long et al., 2023; Chen et al., 2024; Tang et al., 2024c) focuses on generating multi-view images with sufficient 3D consistency to be used for reconstruction. We also adopt this approach of directly generating consistent images for reconstruction. But unlike the above 3D counterparts, there is no large-scale multi-view synchronized video data. Therefore, we opt to combine geometrical consistency priors and video dynamic priors to generate images.
@@
Video generation   Video generation is an active field that has gained increasing popularity. Recent diffusion-based methods have exhibited unprecedented levels of realism, diversity, and controllability. Video LDM (Blattmann et al., 2023b) is the pioneer work to apply the latent diffusion framework (Rombach et al., 2022) to video generation. The subsequent work SVD (Blattmann et al., 2023a) followed its architecture and made effective improvements to the training recipe. W.A.L.T (Gupta et al., 2023) employed a transformer with window attention tailored for spatiotemporal generative modeling to generate high-resolution videos. VDT (Lu et al., 2024) introduced the video diffusion transformer and a spatial-temporal mask mechanism to flexibly capture long-distance spatiotemporal context. The recently introduced SORA (Brooks et al., 2024) demonstrated a remarkable capability to generate long videos with intuitively physical fidelity. Models trained on large-scale video data can generate videos with realistic dynamics. Besides, video diffusion models can also provide effective prior for 3D generators (Chen et al., 2024; Han et al., 2024; Voleti et al., 2024). Therefore, we build our method on this flourishing domain.
@@
4D generation   Animating category-agnostic objects is challenging and has drawn considerable attention from both academia and industry. Unlike 3D generation, 4D generation requires both consistent geometry and realistic dynamics. Recent works on this domain can be categorized based on the input condition. Some of them create 4D content from text or single image. For instance, MAV3D (Singer et al., 2023) first employs SDS in text-to-4D tasks. 4D-fy (Bahmani et al., 2024) hybridizes different diffusion priors during SDS training. But they suffer from extremely slow generation. DreamGaussian4D (Ren et al., 2023) adopts deformable 3D Gaussian (Yang et al., 2024) as the underlying 4D representation and exports mesh for texture refinement, significantly improving efficiency. Another line of work (Jiang et al., 2024b; Wu et al., 2024b; Yin et al., 2023; Pan et al., 2024b) predicts dynamic objects from a single-view video with largely dictated motion. Consistent4D (Jiang et al., 2024b) proposes to use SDS approach for geometry consistency and frame interpolation loss for temporal continuity. 4DGen (Yin et al., 2023) further grounds the 4D content creation with pseudo labels. Meanwhile, L4GM (Ren et al., 2024) directly predicts consistent 3D Gaussians for each frame. And Jiang et al. (2024a) animates existing 3D assets given their multi-view rendering with 4D-SDS. Efficient4D (Pan et al., 2024b) mimics a photogrammetry-based neural volumetric video reconstruction pipeline by directly generating multi-view videos for reconstruction. Some recent works opt for training diffusion models to generate monocular video (Liang et al., 2024) or multi-view videos (Xie et al., 2024) with spatiotemporal consistency for subsequent explicit reconstruction. Compared to previous works, our framework can efficiently generate diverse 4D contents from different input prompts, avoiding the slow and unstable optimization and have potential to continuously benefit from the scalability of underlying diffusion models.
@@ 3 Method
As depicted in figure 2, the proposed 4D generation framework Diffusion2 adopts a two-stage pipeline: dense observation synthesis followed by reconstruction. In this section, we will first elaborate on how to generate dense multi-view multi-frame images for reconstruction through a highly parallelizable denoising process by reconciling the pre-trained video diffusion prior and multi-view diffusion prior while pointing out why it is feasible (stage-1), and then briefly introduce how to robustly reconstruct 4D content from the sampled images (stage-2).
3.1 Estimating scores of the image matrix
@@
In this stage, our goal is to generate highly consistent dense multi-frame multi-view images for reconstruction, which can be denoted as a matrix of images:
@@
Here, $\mathrm{x}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{N}$ is a series of images with $N$ frames or $N$ views, $\nabla_{\mathrm{x}}\log p\left(\mathrm{x};\sigma\right)$ is the score function, which can be estimated as $\nabla_{\mathrm{x}}\log p\left(\mathrm{x};\sigma\right)\approx\left(D_{\theta}(\mathrm{x};\sigma)\right)/\sigma^{2}$ (Karras et al., 2022; Blattmann et al., 2023a), where $D_{\theta}(\mathrm{x};\sigma)$ is a neural network trained via denoising score matching. We want to extend the above formulation to the generation of the whole image matrix, i.e., sampling $\mathcal{I}$ through equation 2 according to its score function $\nabla\log p\left(\mathcal{I};\sigma(t)\right)$. The question is, how do we estimate the score of the joint distribution of $V\times F$ images? Due to the scarcity of available synchronized multi-view videos and the potentially huge memory footprint of simultaneously consuming $V\times F$ images, it is impractical to train a neural network to directly predict the score function of an image matrix with densely distributed views and sufficient frames for substantial motion.
Fortunately, image matrix $\mathcal{I}$ in nature has a nice property, such that we can approximate the $\nabla\log p\left(\mathcal{I};\sigma(t)\right)$ by combining existing video and multi-view score estimators, as claimed in the following theorem:
Theorem 3.1.
For $\mathrm{x}=I_{i,j}$, we have
$\nabla_{\mathrm{x}}\log p({\mathcal{I}};\sigma(t))=\nabla_{\mathrm{x}}\log p({\mathcal{I}}_{\{1:V\},j};\sigma(t))+\nabla_{\mathrm{x}}\log p({\mathcal{I}}_{i,\{1:F\}};\sigma(t))-\nabla_{\mathrm{x}}\log p(I_{i,j};\sigma(t)).$ (3)
Here $\nabla_{\mathrm{x}}\log p({\mathcal{I}}_{i,\{1:F\}})$ and $\nabla_{\mathrm{x}}\log p({\mathcal{I}}_{\{1:V\},j})$ can be predicted by the off-the-shelf video diffusion model and the multi-view diffusion model. This assertion implies that we can sample the desired image matrix by progressively denoising from pure Gaussian noise using the summation of two estimated scores for its row and column, which can be obtained from the pre-trained multi-view and video diffusion models respectively. The derivation of theorem 3.1 is established on our core assumption about the structure of $p(\mathcal{I})$ (refer to appendix A.3 for detailed proof):